# Improved Physics-informed neural networks loss function regularization with a variance-based term

### v2

@@ Abstract
In machine learning and statistical modeling, the mean square or absolute error, is commonly used as an error metric, also called a ”loss function”. While effective in reducing the average error, this approach may fail to address localized outliers, leading to significant inaccuracies in regions with sharp gradients or discontinuities. This issue is particularly evident in physics-informed neural networks (PINNs), where such localized errors are expected and affects the overall solution. To overcome this limitation, we propose a novel loss function that combines the mean and the standard deviation of the chosen error metric. By minimizing this combined loss function, the method ensures a more uniform error distribution and reduces the impact of localized high-error regions. The proposed loss function is easy to implement and tested on problems of varying complexity: 1D Poisson equation, unsteady Burger’s equation, 2D linear elastic solid mechanics, and 2D steady Navier Stokes equations. Results demonstrate improved solution quality and lower maximum error compared to the standard mean-based loss, with minimal impact on computational time.
@@
Machine learning, outlier rejection, physics-informed neural networks, regularization, loss function
@@ 1 Introduction
Deep learning and its applications have been growing rapidly in the past decade or so. The main reasons for that are the availability of large datasets and the ground breaking technology of graphics processing units (GPUs) which have made it possible to train large neural network models with huge datasets.
@@
In the recent years, neural networks were presented to solve partial differential equations (PDE), introducing physics-informed neural networks (PINN) [7]. PINN reduced the need for big datasets and regularize outputs since the physics, governed by PDEs, are used to regularize the predictions. The main idea is using the strong approximation capabilities of neural networks [8] to approximate the solution field of PDEs. The solution is obtained by minimizing a combined loss function which represents the data mismatch (initial/ boundary conditions enforcement and/or sparse data in the domain) and the residual of the PDE over points sampled in the spatiotemporal domain called collocation points. The residual is usually obtained using automatic differentiation [9]. Since, the introduction of PINN, they have been applied in several fields including solid mechanics [10], fluid mechanics [11, 12, 13], flow in porous media [14, 15, 16], composites manufacturing [17, 18], and others.
@@
The article is organized as follows: section 2 is devoted for the introduction of the new proposed loss function, section 3 offers different examples from physics-informed neural networks and their results and comparison to standard loss function. Section 4 finally discusses several aspects such as computational costs or comparison to other strategies, before drawing a conclusion on the proposed loss.
@@ 2.1 Physics-informed neural networks
The residual of the PDE is, then, calculated with automatic differentiation [9]. A combined loss function is formed using both the knowledge of the initial and boundary conditions as well as the physics represented by the residual of the PDE. The loss function is written as:
@@
$\mathcal{L}_{0}=\frac{1}{N_{0}}\sum_{i=1}^{N_{0}}||u(t_{0}^{i},\mathbf{X}_{0}^{i})-u_{0}^{i}||^{2},$ (7)
$\mathcal{L}_{D}=\frac{1}{N_{D}}\sum_{i=1}^{N_{D}}||u(t_{D}^{i},\mathbf{X}_{D}^{i})-u_{D}^{i}||^{2},$ (8)
$\mathcal{L}_{N}=\frac{1}{N_{N}}\sum_{i=1}^{N_{N}}||B(u(t_{N}^{i},\mathbf{X}_{N}^{i}))-f_{N}^{i}||^{2},$ (9)
$\mathcal{L}_{r}=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}||u_{t}+\mathcal{N}(u;\lambda)||_{(t_{r}^{i},\mathbf{X}_{r}^{i})}^{2},$ (10)
@@ 2.2 Proposed loss function
In general, commonly used loss functions take the form of a mean of a chosen error:
@@
The added term to the modified loss function represents the standard deviation of the error function $e$. The full loss function with the added term can be writtne as:
@@
where $\bar{e}$ is the error mean which can be obtained as $\dfrac{\sum_{i=1}^{N}e_{i}}{N}$. $\alpha$ is a hyperparameter that takes values between 00 and $1$ which identifies the contribution of each term in the loss function.
@@ 3 Numerical examples
In all cases, we vary $\alpha$ from 0 to 1 to study the effect of this hyperparameter on the accuracy of the solutions, and we have the same initialization of neural network parameters (with Xavier initialization) each time $\alpha$ is changed to ensure a fair comparison.
@@
$N_{u}$ is the number of training points for the boundary conditions, $N_{f}$ is the number of collocation points used for the PDE residual, $u_{\text{pred}}$ is the predicted solution from the neural network, $u_{\text{data}}$ is the data for boundary conditions, $x_{u}^{i}$ are the spatial coordinates for data points, $x_{f}^{i}$ are the spatial coordinates for collocation points.
The solutions are obtained with 4000 Adam iterations. 100 equally spaced collocation points are used for the training, while the boundary conditions are exactly satisfied by construction. Since an analytical solution to this problem exists, we plot the evolution of the L2 norm of the absolute error between the PINN solution and the analytical solution, vs. the number of iterations for different values of $\alpha$, and the two solutions for the best $\alpha$ value in figure 1.
Figure 1: On the left, the L2 norm of the absolute error vs. the number of iterations for different values of $\alpha$ for the 1D Poisson’s example. On the right, a plot of the analytical solution along with the PINN solution for the best $\alpha$ (0.8).
As shown in figure 1, the best solution is obtained with an $\alpha$ value of 0.8, so adding the variance of the error does improve the accuracy of the prediction. However, the gain in accuracy is not large compared to the classical MSE loss ($\alpha=1$). This can be explained due to the simplicity and the smoothness of the solution. In the next examples, this improvement will be more pronounced.
3.2 unsteady 1D Burger’s problem
We now consider an unsteady example, which represents a standard test case in scientific machine learning, namely the unsteady Burger’s problem in one dimension as introduced by Raissi et al. [7]. The Burger’s equation along with the initial and boundary conditions are written as:
$u_{t}+uu_{x}$ $=(0.01/\pi)u_{xx},$ $x\in[-1,1],\,t\in[0,1],$ (22)
@@
The solutions are obtained using 5000 Adam iterations. 10,000 randomly sampled collocation points are used for training, while the initial and boundary conditions are automatically satisfied by construction. The computational cost using MSE loss is 24.9 seconds and 25.1 seconds when using the variance-based loss. For this case too, an analytical solution exists [30]. First, the L2 norm of the absolute error between the PINN and analytical solutions vs. iterations is plotted for different values of $\alpha$ as shown in figure 2.
Figure 2: L2 norm of the absolute solution error vs iterations for different values of $\alpha$ for the Burger’s problem.
@@
(a) Analytical solution of the Burger’s problem.
(b) Burger’s Solution using MSE loss $\alpha=1$.
(c) Burger’s Solution using new proposed loss $\alpha=0.8$.
@@
The solutions are obtained using 40,000 Adam iterations. 2500 randomly sampled collocation points are used for the training, while the boundary conditions are automatically satisfied by construction. The computational time using MSE loss is 444 seconds vs 450 seconds using the new variance-based loss.
Since an analytical solution of the problem exists, we plot the L2 norm of the absolute error between the predicted and analytical solutions, for both the displacements in the x and y directions, vs. the number of iterations, for different values of $\alpha$ in figure 5.
Figure 5: The L2 norm of the absolute error of x-displacement (left) or y-displacement (right), vs. the number of iterations for different values of $\alpha$ for the 2D linear elasticity example.
@@
(b) Solution of $u_{x}$ using MSE loss $\alpha=1$.
(c) Solution of $u_{x}$ using new proposed loss $\alpha=0.6$.
@@
(b) Solution of $u_{y}$ using MSE loss $\alpha=1$.
(c) Solution of $u_{y}$ using new proposed loss $\alpha=0.6$.
@@
From figures 6 and 7, it is clear that using the newly proposed loss function provides better solutions than using a classical mean squared error. As can be seen in for the displacement in the x direction in figure 6, the maximum absolute error is nearly reduced by a factor of 30. notice that the solution using MSE has a localized error region, while this is not the case for the variance-based loss.
@@ 3.4 Nonlinear case in fluid mechanics
[…]
@@
$\rho\left(u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}\right)=-\frac{\partial p}{\partial y}+\mu\left(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\right),\$ (34)
@@
In this example, 200,000 Adam iterations are employed. The computational time is nearly 55 minutes for MSE loss vs 58 minutes for our proposed loss. 10,176 collocations points are used for training, which are generated from meshing the domain with triangular elements, to be used for the CFD simulation (P1-P1 finite elements with Fenics), as well. The L2 norm of the absolute error between the PINN and CFD solutions is plotted in figure 9 for different $\alpha$ values.
Figure 9: On the left, the L2 norm of the absolute error of the pressure vs. the number of iterations for different values of $\alpha$ for the 2D Navier-Stokes example. On the right, the L2 norm of the absolute error of velocity magnitude vs. the number of iterations for different values of $\alpha$.
@@ 4.1 Computational cost and implementation
4.2 Comparison with Gradient-enhanced PINN
A numerical comparison is performed using Gradient-enhanced PINN (gPINNs) [26] with our proposed loss for the solid mechanics example introduced in section 3.3. The error plots are shown in figure 12.
Figure 12: On the left, the L2 norm of the absolute error of x-displacement vs. the number of iterations for different values of $\alpha$ for the 2D linear elasticity example along with gPINNs solution. On the right, the L2 norm of the absolute error of x-displacement vs. the number of iterations for different values of $\alpha$ along with gPINNs solution.
From the figure, we can see that gPINNs has a similar effect as using our new proposed loss function. It provides a marginal improvement to our solution, however at the expense of a high computational cost. The cost of G-PINN for this problem is more than 3$\times$ that of the original PINN, whereas our proposed modification has negigible computational overhead. Moreover, gPINNs is more difficult to implement, while our proposed regularization is straightforward.
@@ 4.3 Conclusion
In this article, we have proposed a new loss function combining MSE and variance to regularize PINNs. In a variety of examples, the PINN solution clearly improves when applying the new variance-based loss function. This modification effectively reduces the maximum error and minimizes localized error regions, leading to an overall enhancement in solution behavior.
In most of our cases, we noticed that an $\alpha$ of 0.8 provides the best results. This is an empirical finding for which we do not have an explanation yet and needs further testing to be able to generalize to other problems.
Only the Adam optimizer was used in all the cases, in the Burgers and linear elasticity examples, it is enough to obtain good accuracy. For the Navier-Stokes example, Adam did not provide the best solution. In future work, we will explore whether a more powerful optimizer can be used along with Adam such as BFGS to improve the solution. But for the purpose of testing our new loss function, Adam optimizer was enough to demonstrate the usefulness of our proposal.
The added term to the loss function is simple to implement, and the low computational overhead is largely outweighed by the solution improvement.

### v3

@@ Abstract
In machine learning and statistical modeling, the mean square or absolute error is commonly adopted as an error metric, also called a ”loss function”. While effective in reducing the average error, this approach may fail to address localized outliers, leading to significant inaccuracies in regions with sharp gradients or discontinuities. This issue is particularly evident in physics-informed neural networks (PINNs), where such localized errors are expected and affect the overall solution. To overcome this limitation, we propose a novel loss function that combines the mean and the standard deviation of the chosen error metric. By minimizing this combined loss function, the method ensures a more uniform error distribution and reduces the impact of localized high-error regions. The proposed loss function is easy to implement and tested on problems of varying complexity: 1D Poisson equation, unsteady Burger’s equation, 2D linear elastic solid mechanics, and 2D steady Navier Stokes equations. Results demonstrate improved solution quality and lower maximum error compared to the standard mean-based loss, with minimal impact on computational time.
@@
Machine learning; outlier rejection; physics-informed neural networks; regularization; loss function
@@ 1 Introduction
Deep learning and its applications have been growing rapidly in the past decade or so. This is mainly due to the availability of large datasets and the groundbreaking technology of graphics processing units (GPUs) which have made it possible to train large neural network models with huge datasets.
@@
In the recent years, neural networks were presented to solve partial differential equations (PDE), introducing physics-informed neural networks (PINN) [7]. PINN reduce the need for big datasets and regularize outputs since the physics, governed by PDEs, are naturally regularizing the predictions. The main idea is to approximate the solution field of PDEs thanks to the strong approximation capabilities of neural networks [8] . The solution is obtained by minimizing a combined loss function which represents the data mismatch (initial/ boundary conditions enforcement and/or sparse data in the domain) and the residual of the PDE over points sampled in the spatiotemporal domain called collocation points. The residual is usually obtained by automatic differentiation [9]. Since, the introduction of PINN, they have been applied in several fields including solid mechanics [10, 11], fluid mechanics [12, 13, 14, 15], flow in porous media [16, 17, 18], topology optimization [19, 20], composites manufacturing [21, 22], fracture mechanics [23, 24], and many others.
@@
The article is organized as follows: section 2 is devoted to the introduction of the new proposed loss function, section 3 offers different examples from physics-informed neural networks, and their results and comparison to standard loss function. Section 4 finally discusses several aspects such as computational costs or comparison to other well-known strategies, before drawing a conclusion on the proposed loss.
@@ 2.1 Physics-informed neural networks
The residual of the PDE is, then, calculated with automatic differentiation [9]. A combined loss function is formed with both the knowledge of the initial and boundary conditions as well as the physics represented by the residual of the PDE. The loss function is written as:
@@
[…]
@@ 2.2 Proposed loss function
In general, common loss functions take the form of a mean of a chosen error:
@@
The added term to the modified loss function represents the standard deviation of the error function $e$. The full loss function with the added term is written as:
@@
where $\bar{e}$ is the error mean which can be obtained as $\dfrac{\sum_{i=1}^{N}e_{i}}{N}$. $\alpha$ is a hyperparameter that takes values between 0 and 11 which identifies the contribution of each term in the loss function.
@@ 3 Numerical examples
In all cases, we vary $\alpha$ from 0 to 1 to study the effect of this hyperparameter on the accuracy of the solutions, and we have the same initialization of neural network parameters (with Xavier initialization) each time $\alpha$ is changed to ensure a fair comparison. In all cases, NVIDIA RTX 3500 GPU was used for training.
@@
$N_{u}$ is the number of training points for the boundary conditions, $N_{f}$ is the number of collocation points for the PDE residual, $u_{\text{pred}}$ is the predicted solution from the neural network, $u_{\text{data}}$ is the data for boundary conditions, $x_{u}^{i}$ are the spatial coordinates for data points, $x_{f}^{i}$ are the spatial coordinates for collocation points.
The solutions are obtained with 4000 Adam iterations. Training is performed with 100 equally spaced collocation points, while the boundary conditions are exactly satisfied by construction. The computational cost using MSE loss is 8 seconds and 9 seconds when using the variance-based loss. Since an analytical solution to this problem exists, we plot the evolution of the $L_{2}$ norm of the absolute error between the PINN solution and the analytical solution, vs. the number of iterations for different values of $\alpha$, and the two solutions for the best $\alpha$ value in figure 1.
Figure 1: On the left, the $L_{2}$ norm of the absolute error vs. the number of iterations for different values of $\alpha$ for the 1D Poisson example. On the right, a plot of the analytical solution along with the PINN solution for the best $\alpha$ (0.8).
As shown in figure 1, the best solution is obtained with an $\alpha$ value of 0.8: adding the variance of the error does improve the accuracy of the prediction. However, the gain in accuracy is not large compared to the classical MSE loss ($\alpha=1$). This can be explained due to the simplicity and the smoothness of the solution. In the next examples, this improvement will be more pronounced.
3.2 unsteady 1D Burgers’ problem
We now consider an unsteady example, which represents a standard test case in scientific machine learning, namely the unsteady Burgers’ problem in one dimension as introduced by Raissi et al. [7]. The Burgers’ equation along with the initial and boundary conditions are written as:
[…]
@@
The solutions are obtained with 5000 Adam iterations. Training is carried out with 10,000 randomly sampled collocation points, while the initial and boundary conditions are automatically satisfied by construction. The computational cost using MSE loss is 10 seconds and 11 seconds when using the variance-based loss. For this case too, an analytical solution exists [36]. First, the $L_{2}$ norm of the absolute error between the PINN and analytical solutions vs. iterations is plotted for different values of $\alpha$ as shown in figure 2.
Figure 2: $L_{2}$ norm of the absolute solution error vs iterations for different values of $\alpha$ for the Burgers’ problem.
@@
(a) Analytical solution of the Burgers’ problem.
(b) Burgers’ solution using MSE loss with $\alpha=1$.
(c) Burgers’ solution using newly proposed loss with $\alpha=0.8$.
@@
The solutions are obtained with 40,000 Adam iterations. Training involves 2500 randomly sampled collocation points, while the boundary conditions are automatically satisfied by construction. The computational time using MSE loss is 45 seconds vs 47 seconds using the new variance-based loss.
Since an analytical solution of the problem exists, we plot the $L_{2}$ norm of the absolute error between the predicted and analytical solutions, for both the displacements in the x and y directions, vs. the number of iterations, for different values of $\alpha$ in figure 5.
Figure 5: The $L_{2}$ norm of the absolute error of x-displacement (left) and y-displacement (right), vs. the number of iterations for different values of $\alpha$ for the 2D linear elasticity example.
@@
(b) Solution of $u_{x}$ using MSE loss with $\alpha=1$.
(c) Solution of $u_{x}$ using the newly proposed loss with $\alpha=0.6$.
@@
(b) Solution of $u_{y}$ using MSE loss with $\alpha=1$.
(c) Solution of $u_{y}$ using newly proposed loss with $\alpha=0.6$.
@@
From figures 6 and 7, it is clear that using the newly proposed loss function provides better solutions than using a classical mean squared error. As can be seen in for the displacement in the x direction in figure 6, the maximum absolute error is nearly reduced by a factor of 30. Note that the solution using MSE has a localized error region, while this is not the case for the variance-based loss.
@@ 3.4 Nonlinear case in fluid mechanics
Inlet velocity profileP=0(outlet)P=0\ (outlet)No-slip wall
@@
$\rho\left(u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}\right)=-\frac{\partial p}{\partial y}+\mu\left(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\right),\\$ (34)
@@
In this example, 200,000 Adam iterations are employed. The computational time is nearly 15 minutes for MSE loss vs 16 minutes for our proposed loss. Training involves 10,176 collocation points, generated from meshing the domain with triangular elements, which also serve as input for the CFD simulation (P1–P1 finite elements with Fenics). The $L_{2}$ norm of the absolute error between the PINN and CFD solutions is plotted in figure 9 for different $\alpha$ values.
Figure 9: On the left, the $L_{2}$ norm of the absolute error of the pressure vs. the number of iterations for different values of $\alpha$ for the 2D Navier-Stokes example. On the right, the $L_{2}$ norm of the absolute error of velocity magnitude vs. the number of iterations for different values of $\alpha$.
@@ 4.1 Computational cost and implementation
4.2 Comparison with Gradient-enhanced PINN and Huber loss
A numerical comparison is performed using Gradient-enhanced PINN (gPINNs) [32] and Huber loss along with our proposed loss for the solid mechanics example introduced in section 3.3. The error plots are shown in figure 12.
Figure 12: On the left, the $L_{2}$ norm of the absolute error of x-displacement vs. the number of iterations for different values of $\alpha$ for the 2D linear elasticity example along with gPINNs and Huber loss solutions. On the right, the $L_{2}$ norm of the absolute error of x-displacement vs. the number of iterations for different values of $\alpha$ along with gPINNs and Huber loss solutions.
For the Huber loss case, we chose $\delta=1$, where $\delta$ controls the steepness of the Huber curve. From the figure, we can see that the Huber loss has almost no improvement compared with the MSE loss. ($\alpha=1$).
[…]
4.3 Effect on collocation points’ density
In this section, we study the effect of the added term in the loss function on the collocation point’s density needed to reach a comparable accuracy using MSE loss. In figure 13, the $L_{2}$ norm of the error is plotted for two cases of $\alpha$ (1.0, 0.5) vs. the number of collocation points.
Figure 13: On the left, the $L_{2}$ norm of the absolute error of x-displacement vs. the number of collocation points for $\alpha$ values $1.0$ and $0.5$, for the 2D linear elasticity example. On the left, the $L_{2}$ norm of the absolute error of y-displacement vs. the number of collocation points for $\alpha$ values $1.0$ and $0.5$, for the 2D linear elasticity example.
From figure 13, the influence of the additional term in the loss function is evident. In particular, for $\alpha=0.5$, the model achieves a given level of accuracy with substantially fewer collocation points compared to the standard MSE loss ($\alpha=1.0$). This suggests that the modified loss function improves the efficiency of the training process by enforcing the governing equations more effectively, thus reducing the reliance on dense spatial sampling.
Moreover, the improved performance indicates that the added term enhances the network’s ability to capture the underlying physical behavior with fewer constraints. This can be especially beneficial in high-dimensional or computationally expensive problems, where training on a large number of collocation points is impractical or computationally expensive.
@@ 4.4 Conclusion
In this article, we have proposed a new loss function combining MSE and variance to regularize PINNs. In a variety of examples, the PINN solution clearly improves when applying the new variance-based loss function. This modification effectively reduces the maximum error and minimizes localized error regions, leading to an overall enhancement in solution behavior. The new loss function is also more efficient than the MSE loss, requiring less collocation points: this could be a critical point in higher dimension problems. Moreover, it appears better in accuracy or efficiency than two other existing enhanced methods, the Huber loss and gPINNs.
In most of our cases, we empirically found that $\alpha$ = 0.8 often yields the best results: a theoretical justification remains an open question.
Only the Adam optimizer was employed in all the cases: in the Burgers and linear elasticity examples, this optimizer is enough to obtain good accuracy. For the Navier-Stokes example, Adam did not provide the best solution. In future work, we will explore whether a more powerful optimizer can be applied along with Adam such as BFGS to improve the solution. But for the purpose of testing our new loss function, Adam optimizer was enough to demonstrate the usefulness of our proposal.
In conclusion, the added term to the loss function is simple to implement, and the low computational overhead is largely outweighed by the solution improvement.